# Disclosure Avoidance for the 2020 Census Demographic and Housing Characteristics File

### v1

@@ Abstract.
[14] describe the concepts and methods used by the Disclosure Avoidance System (DAS) to produce formally private output in support of the 2020 Census data product releases, with a particular focus on the DAS implementation that was used to create the 2020 Census Redistricting Data (P.L. 94-171) Summary File. In this paper we describe the updates to the DAS that were required to release the Demographic and Housing Characteristics (DHC) File, which provides more granular tables than other data products, such as the Redistricting Data Summary File. We also describe the final configuration parameters used for the production DHC DAS implementation, as well as subsequent experimental data products to facilitate development of tools that provide confidence intervals for confidential 2020 Census tabulations.
Ryan Cumings-Menon, Robert Ashmead, Daniel Kifer,†$\diamond$
Philip Leclerc, Matthew Spence, Pavel Zhuravlev, John M. Abowd
U.S. Census Bureau
Cornell University $\diamond$ Penn State University
@@
This article is © 2023 by author(s) as listed above. The article is licensed under a Creative Commons Attribution (CC BY 4.0) International license (https://creativecommons.org/licenses/by/4.0/legalcode), except where otherwise indicated with respect to particular material included in the article. The article should be attributed to the author(s) identified above.
@@ 1. Introduction
The Census Bureau’s 2020 Disclosure Avoidance System (DAS) was used to produce a formally private file in support of the 2020 Census Redistricting Data (P.L. 94-171) Summary File (hereafter redistricting data); the redistricting data file use case is described in more detail by [14]. The DAS takes the confidential Census Edited File (CEF) as input, which is a set of microdata files with components that include data on individuals, housing units, group quarters (GQs), and households; the output of the DAS is called the Microdata Detailed File (MDF). In the interest of a concise exposition, we refer readers to [14] for more detail on the DAS implementation as well as the concepts in the differential privacy (DP) literature [22] that are used by the DAS. This paper is a companion paper to [14]; we build on this earlier work by providing the updates to the DAS that were required to produce formally private output in support of the Demographic and Housing Characteristics (DHC) File.
@@
After describing the DAS at a high level in the next subsection, the remainder of the paper is organized as follows. Section 2 describes the schemas and constraints used within DHCH and DHCP DAS implementations. Section 3 describes updates to the computational environment used for the DAS DHC implementation. Section 4 describes both how the DAS team went about supporting the choice of production DHC DAS configuration parameters, as well as the production privacy-loss budget (PLB) allocations used to define the noisy measurement files (NMF) for both types of DHC implementations. Section 5 provides summary metrics for the production 2020 DHC file release. One significant advantage of the formally private methods used by the DAS is that they are sufficiently transparent to allow for accurate inferences, i.e., confidence intervals and hypothesis tests, about the underlying CEF-based tabulations. Section 6 concludes with a description of upcoming experimental data products to support the creation of these inferences.
@@ 1.1. A Brief Overview of DAS
Throughout this paper we will make frequent references to [14], but, to make this paper more self-contained, this section outlines the basic steps taken during the DHC and redistricting data. First, after reading the microdata contained in the CEF, these implementations use these confidential microdata to create a flattened fully saturated contingency table, i.e., a flattened histogram, for each census block. In other words, each element (or cell) in the histogram of a given census block corresponds to a possible value of the records, as defined by the product schema, and the value of the cell provides the number of records with this value.
@@
After these histograms have been created in each block geounit, a process called spine optimization is carried out. There are many details related to this process that we will not cover here but that are described by [19]. For the purposes of this paper, it suffices to know that this process defines a hierarchical set of geographic units, which is known as a geographic spine, or simply a spine. For example, the geographic levels for DHC runs, from least to most granular are: US, state, county, prim, tract subset group (TSG), tract subset (TS), optimized block group (OBG), and census block geographic levels. For more detail on the definitions of these geographic levels, see [19]. [^1] State includes state equivalents, namely the District of Columbia. County includes county equivalents. Throughout the paper the geographic unit “block” always means “census block.” The geographic levels used in DHC Puerto Rico runs are the same as the geographic levels listed here, with the exception that a US geographic level is not included. The optimized spine is only used internally within the DAS implementation; the published tabulations do not depend on the definition used for the optimized spine.
@@
After the optimized spine is created, the histogram for each geographic unit (e.g., Fairfax County, Virginia) in the optimized spine is created by summing up the histograms of the blocks it contains. The DAS noisy measurements module, a formally private mechanism that satisfies $\rho-$zero-concentrated differential privacy ($\rho-$zCDP) [16], then forms various marginals from these resulting histograms and adds zero-mean discrete Gaussian noise [17] to these marginals. These noisy marginals are called noisy measurements.
@@
The DAS postprocessing module uses the noisy measurements to create the MDF. Starting from the least granular geographic unit, i.e., the root geographic units, which are either the US or Puerto Rico, it fits a nonnegative and integer histogram to the noisy measurements computed at that geographic level by solving a series of optimization problems, which are described in more detail by [14]. The DAS postprocessing module then fits nonnegative integer histograms to the child geographic units using their noisy measurements, subject to the constraint that their fitted histograms sum to the fitted histogram of the parent geounit. This process is then repeated to define the histogram estimates for each of the remaining geographic levels, from the least to the most granular geographic units. Defining these estimates as solutions to optimization problems also provides a straightforward method to ensure these estimators satisfy a set of constraints, as described by [14]. The specific optimization problems that are used to define the histogram estimate of each geographic unit consist of both least squares and mixed integer programming (MIP) optimization problems. The final step within the DAS implementation is to convert the block geographic level fitted histograms to microdata, which is the output MDF of the DAS implementation.
@@ 2. DHC Schemas and Constraints
In the following subsections we begin by describing the schemas and the constraints of both DHCH and DHCP. Both of these schemas are more granular than their respective redistricting data counterparts. We also ensure that the DHCP MDF is consistent with the MDF of the persons redistricting DAS output and that the DHCH MDF is consistent with the MDF of the units redistricting DAS output. In other words, it is possible to recreate the redistricting persons universe MDF using only the DHCP MDF and to recreate the redistricting units universe MDF using only the DHCH MDF.
@@ 2.1. DHCH Schema
Before providing the schema for the DHCH histogram, we define the DHCH attribute HHTYPE, i.e., household type, as each possible combination of the following attributes.
@@
    SIZE. 7 levels: Indicates if the household includes 1, 2, $\dots$,… , or 7+ members.
@@
The DHCH schema encodes this information in the single attribute HHTYPE because many of the components that make up HHTYPE are related to one another, and encoding the combination of these variables using a single attribute provides a way to exclude combinations of these variables that are impossible in practice. For example, if the SIZE component of HHTYPE indicates the household is composed of only the householder, then SPOUSE must indicate the householder is not married, RELATIVE must indicate the household does not contain a relative of the householder, OWN_CHILD must indicate the householder does not have an own child, and MULTIGEN must indicate the household is not a multigenerational household. In other words, most elements of the cross product of each of these individual components do not actually correspond to a possible level of HHTYPE. Rather than include all 2,240 elements in this naïve cross product of these components of HHTYPE, we decrease the dimension of the resulting DHCH histogram by only including the 522 valid cross product elements.
@@
In addition to the householder histogram, DHCH DAS implementation generates estimates for the detailed cell counts of the units histogram, which provides counts of housing units and GQs rather than counts of householders. The DHCH units histogram has a schema with only one attribute other than GEOCODE, which is known as TENVACGQ (35 levels). TENVACGQ provides the tenure or vacancy status for each housing unit and the detailed GQ type category for each GQ. Specifically, the first four levels of TENVACGQ correspond to levels of occupied housing units; these levels indicate if a property is mortgaged, owned, rented, or occupied by someone who is not the owner or a renter. The subsequent seven levels correspond to types of vacant housing units; these levels indicate if the housing unit is vacant and for rent, vacant and rented but not occupied, vacant and for sale, vacant and sold but not occupied, vacant because the housing unit is for seasonal, recreational, or occasional use, vacant because the property is for migrant workers, or vacant for other reasons. The subsequent 24 levels correspond to the detailed GQ type categories used in DHCH. The units histogram is estimated at the same time as the householder histogram during the DHCH DAS implementation because a constraint links both of these histograms to one another. This constraint, and all other constraints in the 2020 DHCH production implementation settings, are provided in the list below.
  1. (C1)
@@
  3. (C3)
    Structural zero constraints are included to ensure consistency of attributes with the SIZE component of HHTYPE. For example, households of size one cannot have a householder with age less than 60 and have a household member greater than 60.
  4. (C4)

### v2

@@ Abstract.
[abowd20222020] describe the concepts and methods used by the Disclosure Avoidance System (DAS) to produce formally private output in support of the 2020 Census statistical data product releases, with a particular focus on the DAS implementation that was used to create the 2020 Census Redistricting Data (P.L. 94-171) Summary File. In this paper we describe the updates to the DAS that were required to release the Demographic and Housing Characteristics (DHC) File, which provides more granular tables than other statistical data products, such as the Redistricting Data Summary File. We also describe the final configuration parameters used for the 2020 production DHC DAS implementation, error metrics for these production statistical data products, and plans for future experimental data products that provide confidence intervals for confidential 2020 Census tabulations.
[…]
@@
1. Media Summary
One of the more granular statistical data products produced for the 2020 Census is the Demographic and Housing Characteristics (DHC) File. This data product provides tabulations for detailed demographic groups all the way down to census blocks, which are the smallest set of geographic areas used by the Census Bureau. This granularity makes this data product particularly useful to researchers, governmental entities, and businesses.
Like the other 2020 Census statistical data products, the Census Bureau’s Disclosure Avoidance System (DAS) was used on the DHC tabulations to protect the confidentiality of respondents. The DAS tracks the cumulative privacy-loss of respondents due to the publication of the 2020 Census data products using a formally private accounting framework. A variety of implementation and tuning choices were required for the DHC DAS execution to balance multiple trade-offs, including between utility for various DHC use cases, confidentiality of respondents, and the granularity of tabulations. To make these choices, the Census Bureau released two DHC demonstration data products based on 2010 data and collected feedback on the feasibility of potential 2020 DHC use cases from subject-matter experts (SMEs), both internally to the Census Bureau and externally. Final approval of these settings and algorithmic choices was made by the Data Stewardship Executive Policy (DSEP) Committee.
One advantage of this approach relative to the disclosure avoidance methods used in previous decennial censuses is transparency; the final DHC DAS tuning parameter choices, DHC DAS codebase, the intermediate noisy tabulation measurements generated within DHC DAS production execution, and approximate accuracy metrics of the perturbations introduced by the DHC DAS are now publicly available. This transparency also makes it possible for the Census Bureau to formulate confidence intervals that account for the perturbations introduced by the 2020 DAS.
@@ 2. Introduction
The Census Bureau’s 2020 Disclosure Avoidance System (DAS) was used to produce a formally private file in support of the 2020 Census Redistricting Data (P.L. 94-171) Summary File (hereafter redistricting data); the redistricting data file use case is described in more detail by [abowd20222020]. The DAS takes the confidential Census Edited File (CEF) as input, which is a set of microdata files with components that include data on individuals, housing units, group quarters (GQs), and households; the output of the DAS is called the Microdata Detailed File (MDF). In the interest of a concise exposition, we refer readers to [abowd20222020] for more detail on the DAS implementation as well as the concepts in the differential privacy (DP) literature [dwork2006calibrating] that are used by the DAS. This paper is a companion paper to [abowd20222020]; we build on this earlier work by providing the updates to the DAS that were required to produce formally private output in support of the Demographic and Housing Characteristics (DHC) File.
@@
In addition to providing the algorithmic changes and production settings used for the 2020 DHC DAS production executions, this paper outlines the process the Census Bureau used to investigate candidate DHC DAS production settings, i.e., the DHC accuracy experiments. These accuracy experiments involved collecting, and attempting to satisfy, accuracy targets for DHC statistical data product use cases with minimal privacy-loss budget (PLB). These accuracy targets were provided by both internal Census subject-matter experts (SMEs) and external SMEs, including participants of academic conferences and feedback provided through a National Academies of Science, Engineering, and Medicine’s Committee on National Statistics (CNSTAT) meeting in June of 2022 [cnstat]. External SMEs based their feedback on two 2020 DHC demonstration data product DAS executions that used 2010 data as input, as is described in more detail in Section 5. Accuracy metrics from these experiments, and input from SMEs, were used by the Data Stewardship Executive Policy Committee (DSEP) to set the final 2020 DHC DAS production settings.
The basic top-down structure used in both the 2020 Redistricting Data (P.L. 94-191) Summary File and the Demographic and Housing Characteristics file was an algorithmic decision primarily motivated by the publication requirements of redistricting, specifically, the necessity to publish data with census blocks as the geographic atom. The executive decision that all DHC tables must be aggregation-consistent with the redistricting tables implied that census blocks remain the geographic atom for DHC. Unlike the redistricting data, which were used to draw voting districts whose boundaries could not be specified in advance, all DHC tables used geographic boundaries that were predetermined in the 2020 Census tabulation geography. This allowed the algorithmic and SMEs to tune the geographic spine used in DHC to improve accuracy with smaller allocations of PLB. These efforts, described in this paper, illustrate the intellectual contribution of the DHC for publishing very high dimensional table sets. Other practitioners can use similar methods.
One significant advantage of the formally private methods used by the DAS is that they are sufficiently transparent to allow for accurate statistical inferences, i.e., confidence intervals and hypothesis tests, of the underlying confidential CEF-based tabulations. This paper also describes areas of ongoing work aimed at leveraging this advantage to provide experimental data products with confidence intervals for CEF-based tabulations.
After describing the DAS at a high level in the next subsection, the remainder of the paper is organized as follows. Section 3 describes the schemas and constraints used within DHCH and DHCP DAS implementations. Section 4 describes updates to the computational environment used for the DAS DHC implementation. Section 5 describes the process used to set the 2020 production DHC DAS configuration parameters, including the production PLB allocations that define the distributions of the noisy measurements observed in the DAS DHC implementations. Section 6 provides summary metrics for the production 2020 DHC file release. Section 7 concludes with a description of upcoming experimental data products to support statistical inferences on CEF-based tabulations.
@@ 2.1. A Brief Overview of DAS
Throughout this paper we make frequent references to [abowd20222020], but, to make this paper more self-contained, this section outlines the basic steps taken during the DHC and redistricting data DAS executions. First, after reading the microdata contained in the CEF, these implementations use these confidential microdata to create a flattened fully saturated contingency table, i.e., a histogram, for each census block. In other words, each element (or cell) in the histogram of a given census block corresponds to a possible value of the records, as defined by the product schema, and the value of the cell provides the number of records with this value. Since microdata and histogram data formats are referenced throughout this paper, the next example describes these formats and how one can map between these two encodings of the data.
Table 1. An example dataset in microdata format
STATE_GEOCODE HISPANIC SEX
24 0 0
24 1 0
24 1 1
55 0 1
55 0 1
Example 1.
In this example we describe how to map the dataset given in Table 1 from microdata format to a histogram. We use microdata to mean the data format in which each row provides an observation and each column provides the attribute values of the observations. We refer to the list of attributes and the universe of possible levels (or values) for each attribute as the schema of the data. For example, the dataset in Table 1 has the following schema.
  1. (1)
    STATE_GEOCODE. 51 levels: Provides the FIPS state geographic level identification code. Note that the District of Columbia is included as a state equivalent in this geographic level.
  2. (2)
  3. (3)
    SEX. 2 levels: Indicates if the respondent is Male or Female.
In contrast, a histogram provides the count of records with each possible combination of attribute levels. Since the DAS implementation uses histograms of individual geographic units, next we describe how to convert this dataset to the histogram encoding for the state with a STATE_GEOCODE of 24. This can be done by first filtering the microdata dataset given in Table 1 to the three rows with STATE_GEOCODE=24.=24. Second, we sequentially consider each possible combination of the levels of the remaining attributes in lexicographic order, and collect the count of records with the corresponding combination of levels into a vector, which provides the final histogram for this geographic unit. For example, the first element of this lexicographic ordering on the levels of the HISPANIC and SEX attributes are cells with HISPANIC=0 and SEX=0. Since there is one record with these levels, the first such record count is 1. After continuing this process for the remaining level combinations, we can see that the histogram for this geographic unit is given by $\boldsymbol{x}=(1,0,1,1)^{\top}.$
Histograms are a convenient format to use within the DAS implementation because they allow for many operations to be specified with linear algebra. For example, if $A\otimes B$ denotes the Kronecker product of the matrices $A$ and $B,$ we can use the histogram $\boldsymbol{x}$ to compute the SEX marginal for this geographic unit using
$\left(\begin{bmatrix}1&1\end{bmatrix}\otimes\begin{bmatrix}1&0\\ 0&1\end{bmatrix}\right)\boldsymbol{x}=\begin{bmatrix}1&0&1&0\\ 0&1&0&1\end{bmatrix}\begin{bmatrix}1\\ 0\\ 1\\ 1\end{bmatrix}=\begin{bmatrix}2\\ 1\end{bmatrix}$.
This feature is particularly important when specifying linear constraints in optimization models, in which case the histogram vectors are variables that are being optimized over rather than fixed constants, since all constraints used within DAS other than some integrality constraints can be specified as linear constraints. ∎
@@
After these histograms have been created in each block geographic unit, a process called spine optimization is carried out. There are many details related to this process that we do not cover here but that are described by [cumings2022geographic]. For the purposes of this paper, it suffices to know that this process defines a hierarchical set of geographic units, which is known as a geographic spine, or simply a spine. For example, the geographic levels for the DHC DAS implementation, from least to most granular are: US, state, county, prim, tract subset group (TSG), tract subset (TS), optimized block group (OBG), and census block geographic levels. For more detail on the definitions of these geographic levels, see [cumings2022geographic]. [^1] State includes state equivalents, namely the District of Columbia. County includes county equivalents. Throughout the paper the geographic unit “block” always means “census block.” The geographic levels used in DHC Puerto Rico executions are the same as the geographic levels listed here, with the exception that a US geographic level is not included. Prim geographic units are defined as the most granular geography unit used internally by the Population Estimates and Projections (PEP) Area of the Census Bureau. This geographic level includes entities defined as the intersections of counties and other geographic units, such as functioning places and minor civil divisions. The DHC DAS executions used different spines than the ones used for the redistricting data file DAS executions for two reasons. First, the maximum number of child geographic units of any parent geographic unit is lower for the DHC optimized spine; this is required to ensure that the optimization solves used within DAS executions, which are described below in more detail, are computationally feasible. Second, the differences in use cases between the DHC and redistricting data file releases led to updates to the off-spine geographic entities that the spine optimization routines target accuracy in. For example, one use case for the DHC tabulations is forecasting school enrollment figures, so school districts were added to this set of off-spine entities that are targeted for accuracy in the DHC DAS implementation. Even though the geographic spine used internally within the DAS implementation does not correspond to the Census Bureau Geography Division’s standard tabulation spine, the published 2020 DHC tabulations are still counts for geographic entities on the standard tabulation spine.
@@
After the optimized spine is created, the histogram for each geographic unit (e.g., Fairfax County, Virginia) in the optimized spine is created by summing up the histograms of the blocks it contains. The DAS noisy measurements module, a formally private mechanism that satisfies $\rho$-zero-concentrated differential privacy ($\rho$-zCDP) [bun2016concentrated], then forms various marginals from these resulting histograms and adds zero-mean discrete Gaussian noise [canonne2020discrete] to these marginals. These noisy marginals are called noisy measurements.
@@
The DAS postprocessing module uses the noisy measurements to create the MDF. Starting from the least granular geographic unit, i.e., the root geographic units, which are either the US or Puerto Rico, it fits a nonnegative and integer histogram to the noisy measurements computed at that geographic level by solving a series of optimization problems, which are described in more detail by [abowd20222020]. The DAS postprocessing module then fits nonnegative integer histograms to the child geographic units using their noisy measurements, subject to the constraint that their fitted histograms sum to the fitted histogram of the parent geographic unit. This process is then repeated to define the histogram estimates for each of the remaining geographic levels, from the least to the most granular geographic units. Defining these estimates as solutions to optimization problems also provides a straightforward method to ensure these estimators satisfy a set of constraints, as described by [abowd20222020]. The specific optimization problems that are used to define the histogram estimate of each geographic unit are carried out in a series of optimization passes, each consisting of one nonnegative least squares (NNLS) solve to estimate a vector of non-negative count estimates and one mixed integer programming (MIP) solve to convert these estimates to integers. These optimization passes are repeated to estimate and fix nonnegative and integer counts of progressively more granular tabulations, and the final optimization pass estimates and fixes the detailed histogram counts of the geographic unit. The final step within the DAS implementation is to convert the block geographic level fitted histograms to microdata, which is the output MDF of the DAS implementation.
After the production DAS executions were complete, the 2020 Census tabulation tables were created using the output MDFs. Part of our planning process for the 2020 Census statistical data product publications involved updates to the tables that were released as part of the 2010 Census statistical data products. This included adding, removing, and altering the granularity of the published tables, as well as moving some tables to the more specialized Group II statistical data products, which are the 2020 Census Detailed DHC-A, Detailed DHC-B, and the S-DHC statistical data products. In aggregate, these updates resulted in a reduction in the total number of tables published as part of all statistical data products released for the 2020 Census. The updates that were most relevant to our adoption of the DHC DAS implementation for the 2020 Census were the removal of the tables included in the 2020 Census Group II statistical data products from the DHC publications because supporting tabulations that included the detailed race and ethnicity categories considered in these Group II products (e.g., German, Chinese, Mexican, etc., as opposed to the less granular race and ethnicity categories used in DHCH/DHCP schemas like White, Asian, Hispanic/Latino, etc.) would have significantly increased the dimensions of the optimization models in the DHC DAS implementation. [tableCrosswalk] provides a description of the tables used for the 2020 Census publications, along with a crosswalk between the 2010 Census and the 2020 Census published tables.
@@ 3. DHC Schemas and Constraints
In the following subsections we begin by describing the schemas and the constraints of both DHCH and DHCP. Both of these schemas are more granular than their respective redistricting data counterparts. We also ensure that the DHCP MDF is consistent with the MDF of the persons redistricting DAS output and that the DHCH MDF is consistent with the MDF of the units redistricting DAS output. In other words, it is possible to recreate the redistricting persons universe MDF using only the DHCP MDF and to recreate the redistricting units universe MDF using only the DHCH MDF. Constraints that ensure this consistency between these statistical data products are imposed for two reasons. First, it avoids the possibility of confusion caused by conflicts between these two data products. Second, these constraints allow us to ensure accuracy of the less-granular queries that are fixed by the redistricting data file DAS evaluations without expending additional PLB. Constraining to the redistricting data file production DAS execution outputs also ensures that the state total population invariant constraints that were imposed in these DAS executions hold for tabulations based on the DAS DHC implementation.
The CEF is created using edit and imputation methods to transform the unedited data collected during the decennial Census to data that satisfies certain conditions, called edit constraints, and the remaining constraints imposed in DHC DAS executions described in this section are taken directly from these edit constraints. For example, one such edit constraint ensures that the age of at least one person in each household is at least 15.
@@ 3.1. DHCH Schema
As described in Section 2, the published DHCH tabulations provide counts for two distinct categories of entities: households and of housing units/GQs. Note that the structure in which a household resides is called a housing unit, and people residing in GQs are not viewed as household members. The majority of the published DHCH tabulations are in the first of these two categories, i.e., they provide counts of households. Since each household has exactly one householder, these counts can alternatively be interpreted as tabulations of householders in certain types of households. While these two categories of tabulations are distinct from one another, they are also related in a few ways. For example, the number of housing units that are owned by their residents with a mortgage in a geographic unit must be equal to the number of households that own their home with a mortgage. These relationships between these two categories of counts provide a way for the noisy measurements of counts of housing units to be used to enhance the precision of count estimates for households. In order to do this, the DHCH DAS implementation estimates counts of these two categories within a single DAS execution by estimating and fixing a secondary units histogram, in addition to the main household histogram. This section describes the schemas of these two histograms, starting with the main household histogram. To do so, we first define one attribute of this main histogram, called HHTYPE, i.e., household type. Each level of HHTYPE is defined as a combination of the following attributes that is actually feasible in practice, which we describe below in more detail.
@@
    SIZE. 7 levels: Indicates if the household includes 1, 2, $\dots,$ or 7+ members.
@@
The DHCH main histogram schema encodes this information in the single attribute HHTYPE because many of the components that make up HHTYPE are related to one another, and encoding the combination of these variables using a single attribute provides a way to exclude combinations of these variables that are impossible in practice. For example, if the SIZE component of HHTYPE indicates the household is composed of only the householder, then SPOUSE must indicate the householder does not live with a spouse or partner, RELATIVE must indicate the household does not contain a relative of the householder, OWN_CHILD must indicate the householder does not have an own child, and MULTIGEN must indicate the household is not a multigenerational household. In other words, most elements of the cross product of each of these individual components do not actually correspond to a possible level of HHTYPE. Rather than include all 2,240 elements in this naïve cross product of these components of HHTYPE, we decrease the dimension of the resulting DHCH histogram by only including the 522 valid cross product elements.
@@
The DHCH units histogram has a schema with only one attribute other than GEOCODE, which is known as TENVACGQ (35 levels). TENVACGQ provides the tenure or vacancy status for each housing unit and the detailed GQ type category for each GQ. Specifically, the first four levels of TENVACGQ correspond to levels of occupied housing units; these levels indicate if a property is mortgaged, owned, rented, or occupied without payment of rent. The subsequent seven levels correspond to types of vacant housing units; these levels indicate if the housing unit is vacant and for rent, vacant and rented but not occupied, vacant and for sale, vacant and sold but not occupied, vacant because the housing unit is for seasonal, recreational, or occasional use, vacant because the property is for migrant workers, or vacant for other reasons. The subsequent 24 levels correspond to the detailed GQ type categories used in DHCH. The units histogram is estimated at the same time as the householder histogram during the DHCH DAS implementation because a constraint links both of these histograms to one another. This constraint, and all other constraints in the 2020 DHCH production implementation settings, are provided in the list below.
  1. (H1)
@@
  3. (H3)
    Structural zero constraints are included to ensure consistency of attributes with the SIZE component of HHTYPE. For example, households of size one cannot have a householder with age less than 60 and have a household member with age greater than 60.
  4. (H4)